# Generalized Nash Equilibrium Solutions in Dynamic Games With Shared Constraints

### v2

@@ 7 Numerical Two-Car Racing Problem
Solving the above KKT equations of any value of $\alpha\in(0,\infty)$ gives a valid GNE solution. The solution set of 2 examples are presented here. Each example contains the calculated GNE solutions of the race for a given initial states of the car on different tracks. The first example is calculated on a straight track. Car 1 is ahead of car 2 but has a velocity disadvantage. Figure 2 shows the solution of the race for different $\alpha$. The solution for $\alpha\rightarrow 0$ gives a straight trajectory (red line) for car 1 and this corresponds to car 1 being more aggressive than car 2. The solution for $\alpha\rightarrow\infty$ gives a straight trajectory (magenta line) for car 2 and this corresponds to car 2 being more aggressive than car 1. The GNE solution changes continuously as $\alpha$ changes. The lower graph shows the cost function of both cars as function of $\alpha$, along with their sum. It can be seen that for $\alpha\approx 2.5$ the sum of cost is minimized, and it corresponds to a trajectory that is very similar to a trajectory in which $\alpha\rightarrow\infty$. On the left side of the graph the velocity and steering profile of the cars are presented.
@@
Figure 2: Racing Problem Solutions on a Straight Track
@@
The second example is calculated on a curved track. Car 1 is ahead of car 2 but has a velocity disadvantage. Figure 3 shows the solution of the race for different $\alpha$. In this example there are 2 types of possibles solutions. For $\alpha\rightarrow 0$ Car 2 is much more aggressive and cuts into the corner which enforces Car 1 to take the wider turn. On the other hand, for $\alpha\rightarrow\infty$ car 1 is much more aggressive and it cuts into the corner before car 2 can do it - this enforces car 2 to take the wider turn. The GNE solution doesn’t changes continuously as $\alpha$ changes. The lower graph shows the cost function of both cars as function of $\alpha$ and the their sum. It can be seen that for $\alpha\approx 10$ there is a jump in the costs which corresponds to the jump in the solution types. In this example the sum of cost is minimized for an $\alpha$ close to the switching point. On the left side of the graph the velocity and steering profile of the cars are presented.
@@
Figure 3: Racing Problem Solutions on a Curved Track
8 Conclusion
@@
This paper introduces a novel method for extending existing approaches used to solve dynamic games (DG) through Variational Inequality (VI) and Mixed Complementarity Problem (MCP) tools. While these traditional methods have been predominantly restricted to computing the normalized solution, the proposed framework allows for the computation of non-normalized solutions, significantly expanding the solution space in DGs while using the same powerful methods. The proposed method not only identifies non-normalized solutions but also provides a systematic approach to select a Generalized Nash Equilibrium (GNE) from the potentially large set of solutions. By leveraging bi-level optimization, the method enables solution selection based on specific objectives, such as minimizing the aggregate cost, optimizing the performance of individual players, or prioritizing game-specific properties. Furthermore, the examples demonstrate the effectiveness of the proposed framework in practical applications, highlighting its ability to identify diverse GNE solutions and to tailor these solutions to specific objectives. This approach addresses the limitations of normalized solutions by accounting for terms that are neglected in the standard solution. In conclusion, this work expands the computational toolbox for DGs by enabling exploration beyond normalized solutions and by offering a structured methodology for equilibrium selection. These contributions pave the way for more robust and flexible applications of game-theoretic principles in complex, real-world scenarios.

### v3

@@ VII-A Quantitative Analysis of Tuning Aggressiveness
The first example is calculated on a straight track. Car 1 (blue) is ahead of car 2 (green) but has a velocity disadvantage. Figure 2 shows the solution of the race for different $\alpha$. The solution for $\alpha\rightarrow 0$ gives a straight trajectory (red line) for car 2 and this corresponds to car 2 being more aggressive than car 1. The solution for $\alpha\rightarrow\infty$ gives a straight trajectory (cyan line) for car 1 and this corresponds to car 1 being more aggressive than car 2. The GNE solution changes continuously as $\alpha$ changes. The lower graph shows the cost function of both cars as function of $\alpha$. A car can improve its cost by playing more aggressively. On the left side of the graph the velocity and steering profile of the cars are presented. The black line represent the normalized solution.
@@
Figure 2: First Qualitative Example: Racing Problem Solutions on a Straight Track
@@
The second example is calculated on a curved track. Car 1 is ahead of car 2 but has a velocity disadvantage. Figure 3 shows the solution of the race for different $\alpha$. In this example there are 2 types of possibles solutions. For $\alpha\rightarrow 0$ Car 2 is much more aggressive and cuts into the corner which enforces Car 1 to take the wider turn. On the other hand, for $\alpha\rightarrow\infty$ car 1 is much more aggressive and it cuts into the corner before car 2 can do it - this enforces car 2 to take the wider turn.
The GNE solution does not change continuously as $\alpha$ changes. The lower graph shows the cost function of both cars as function of $\alpha$. This multi modal solution was not possible if only the normalized solution is calculated. It can be seen that for $\alpha\approx 10$ there is a jump in the costs which corresponds to the jump in the solution types. On the left side of the graph the velocity and steering profile of the cars are presented.
@@
Figure 3: Second Qualitative Example: Racing Problem Solutions on a Curved Track
VII-B Quantitative Results: A Numerical Study
The examples above show the set of options a race car has at different situations of the race. Choosing the right strategy from all possible is essential to get an advantage. The intuition suggests that the more aggressive a car will act (meaning, choosing lower $\alpha$) the better its performance will be. To prove this, we present a numerical study of a racing simulation of 2 racing cars. Both cars race by solving a racing DG problem as presented in (47).
The scenario investigated is on an L-shaped track where the cars race in the counter-clock direction. The initial state of the cars is randomly initialized along the track with velocities such that the trailing car (called ego) has a velocity advantage over the forward car (Opponent).
Two scenarios are investigated. In the first, both cars implement a strategy that is the result of the normalized solution ($\alpha=1$). In the second scenario, the opponent continues to implement the normalized strategy (meaning, it solve the DG while assuming both car use a normalized solution) while the ego car solves the DG with non-normalized solution ($\alpha=0.05$).
In both cases, the goal of the ego car is to overtake the opponent during a $2[sec]$ race started from the randomized initial conditions. The simulation is closed-loop, meaning that each player solves the DG at each time step, implements a single control input and solves the DG again to get the next time step.
Figure 4: Results of 100 Monte Carlo simulations comparing two different strategies on an L-shaped track
Ego Plays Normalized Ego Plays Non-Normalized
Win Percentage 43% 54%
TABLE I: MC simulation results summary comparing both strategies
The ego has a top speed of $3[m/s]$ while the opponent has a top speed of $2.85[m/s]$, otherwise both cars have the same dynamic model and constraints. The initial state of the cars is randomized with properties presented in the following table:
Description Range
Opponent initial position $[0,L]$
Ego initial position relative to opponent [1.75,1.5][-1.75,-1.5]
Opponent initial velocity $[1.0,2.0]$ m/s
Ego initial velocity relative to opponent $[0.25,0.75]$ m/s
Ego lateral offset [H/3,H/3][-H/3,H/3] m
Opponent lateral offset (relative to ego) [H/8,H/8][-H/8,H/8] m
TABLE II: Randomized Parameters in the Monte Carlo Simulation. $L$ is the track length and $H$ is the track half width.
Table I summarizes the results of this study. When both cars use the normalized strategy (exhibiting similar levels of aggressiveness), the ego car overtakes the opponent in $43\%$ of the scenarios. However, when the ego car adopts a more aggressive strategy than the opponent, it wins $54\%$ of the scenarios—an $11\%$ increase in win probability. Note that the advantage of the non-normalized strategy may be even greater, considering that certain randomized initial conditions may inherently prevent the ego car from winning, regardless of its strategy. . Another visualization of these results is available in Figure 4. Each point on the track represents the initial condition of the ego car in a single racing scenario. A green scenario indicates a win for the ego car (successful overtaking of the opponent), while a red scenario indicates a loss (failure to overtake the opponent). This shows the effectiveness of the non-normalized strategy relative to the normalized one.
Overall, these results clearly demonstrate that a more aggressive strategy increases the likelihood of winning. While this conclusion is intuitive, proving it rigorously is nontrivial. The non-normalized solutions presented in this paper provide a systematic method for controlling the relative aggressiveness of players in a game and analyzing its impact on the final outcome.
Also, while the increase in the win percentage from 43% to 54% may not seem substantial, the theoretical upper bound on the optimal win rate in this setting is difficult to compute. We conjecture that, given the constraints of this scenario, the optimal win percentage when both players pursue non-normalized solutions is likely not significantly higher than 54%.
Clearly, in real-time applications, an identification of the opponent’s aggressiveness needs to be carried out and adjusted dynamically to optimize individual performance while avoiding crashes. Understanding the effects of time-varying opponent aggressiveness is an interesting direction for future research but is beyond the scope of this paper.
VIII Conclusion
@@
This paper introduces a novel method for extending existing approaches used to solve dynamic games (DG) through Mixed Complementarity Problem (MCP) tools. This traditional method has been restricted to computing the normalized solution. The proposed framework allows for the computation of non-normalized solutions, significantly expanding the solution space in DGs while using the same numerical tools. By leveraging bi-level optimization, the method enables solution selection based on specific objectives, such as minimizing the aggregate cost, optimizing the performance of individual players, or prioritizing game-specific properties. Furthermore, it was shown that by interpreting the different GNE solution by the level of relative aggressiveness of each player, a player can get an advantage by playing more aggressively. The examples demonstrate the effectiveness of the proposed framework in practical applications, highlighting its ability to identify diverse GNE solutions and to tailor these solutions to specific objectives. This approach addresses the limitations of normalized solutions by accounting for terms that are neglected in the standard solution. In conclusion, this work expands the computational toolbox for DGs by enabling exploration beyond normalized solutions and by offering a structured methodology for equilibrium selection. These contributions pave the way for more robust and flexible applications of game-theoretic principles in complex, real-world scenarios.
[Uncaptioned image] Mark Pustilnik Mark Pustilnik received his B.S.c in 2007 and M.Sc. in 2013 in Aerospace Engineering from the Technion – Israel Institute of Technology. He is currently a Ph.D. candidate in Mechanical Engineering at the University of California, Berkeley. His research focuses on dynamic games, generalized Nash equilibria, optimal control, and trajectory optimization.
[Uncaptioned image] Prof. Francesco Borrelli Francesco Borrelli received the Laurea degree in computer science engineering from the University of Naples Federico II, Naples, Italy, in 1998, and the Ph.D. degree from ETH-Zurich, Zurich, Switzerland, in 2002. He is currently an Associate Professor with the Department of Mechanical Engineering, University of California, Berkeley, CA, USA. He is the author of more than 100 publications in the field of pre- dictive control and author of the book Constrained Optimal Control of Linear and Hybrid Systems (Springer-Verlag). His research interests include constrained optimal control, model predictive control and its application to advanced automotive control and energy efficient building operation. Dr. Borrelli received the 2009 National Science Foundation CAREER Award and the 2012 IEEE Control System Technology Award. In 2008, he became the Chair of the IEEE Technical Committee on Automotive Control.